# Nonparametric Estimation of Matching Efficiency and Mismatch in Labor Markets via Public Employment Security Offices in Japan, 1972-2024

### v3

@@ Abstract
Keywords: matching efficiency, mismatch
@@ 1.1 Related Literature
This paper contributes to the three strands of literature. First, I examine the trend of matching efficiency in Japanese labor markets via Hello Work nonparametrically using a novel approach (Lange and Papageorgiou 2020), which shows how to nonparametrically identify the matching function and estimate the matching function allowing for unobserved matching efficacy, without imposing the usual independence assumption between matching efficiency and search on either side of the labor market, allowing for multiple types of jobseekers. Lange and Papageorgiou (2020) highlight positive correlations between efficiency and market structure such as tightness and so on, which induces a positive bias in the estimates of the vacancy elasticity whenever unobserved matching efficacy is not controlled for, as is the case in the traditional Cobb-Douglas matching function with constant elasticity parameters. [^1] Brancaccio et al. (2020a, 2023) apply the method to estimate matching function in exporter-ship transactions in the global bulk shipping market. Brancaccio et al. (2020b) summarize practical issues. Implementing their approach, I add updated results from the more flexible approach to the existing findings reported in Kano and Ohta (2005), Kambayashi and Ueno (2006), Sasaki (2008), and Higashi (2018) using the traditional Cobb-Douglas matching function with geographical and occupational category fixed effects to capture geographical and occupational heterogeneity. Table 1 summarizes the previous findings. My findings are also useful for comparison with other countries’ results reported in Bernstein et al. (2022) and Petrongolo and Pissarides (2001).
@@ 2.1 Nonparametric aggregate matching function
I identify the matching function as well as unobserved, time-varying matching efficiency, $A.$ First, I assume that $V$ and $A$ are independent conditional on $U$, that is, $A\perp V\mid U$. Second, I assume that the matching function $m(AU,V):\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}$ has constant returns to scale (CRS). [^5] To follow the original model of Matzkin (2003), we can rewrite $H=m(AU,V)$ into $H/U=m(A,V/U)$ under CRS. Here, $H/U$ and $V/U$ are known as a job-finding rate and market tightness. Then, by applying nonparametric identification results of Matzkin (2003), Proposition 1 of Lange and Papageorgiou (2020) proves that $G(H,U,V)$ identifies $F(A,U)$ and $m(AU,V):\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}_{+}$ up to a normalization of $A$ at one point denoted as $A_{0}$ of the support of $(A,U,V)$. [^6] In Appendix, I report finite sample performance and methodological extension with Monte Carlo simulation. Based on simulation results with sample size $T=50$, the sample size in this paper is enough to recover matching efficiency well. The code is available on the author’s Github. Also, see Brancaccio et al. (2020b) for practical issues. The approach is used to estimate a matching function in a trade model (Brancaccio et al. 2020a, 2023).
@@ 2.2 Nonparametric mismatch
I construct a nonparametric version of a mismatch index which is a measure of the fraction of hires foregone due to mismatch. Suppose there are $G$ local markets in time $t$ with a given number of vacancies, unemployed workers, and aggregate matching efficiencies in the economy. The number of hires $H_{gt}$ in market $g$ in time $t$ is determined by a matching function $m(A_{gt}U_{gt},V_{gt})$. Note that, unlike Şahin et al. (2014), I allow any specification of $m$ satisfying CRS and independence of $A_{gt}$ and $U_{gt}$ conditional on $V_{gt}$. Given matching efficiency $A_{gt}$ and the number of vacancies $V_{gt}$, a social planner maximizes the total hires in the economy by distributing a given number of unemployed workers to each labor market.
@@
$\frac{\partial m}{\partial U}\left(A_{1t}U_{1t}^{*},V_{1t}\right)=\cdots=\frac{\partial m}{\partial U}\left(A_{Gt}U_{Gt}^{*},V_{Gt}\right)$ (1)
where * denotes the planner’s allocation. The equilibrium condition means that keeping total unemployed workers $U_{t}=\sum_{g=1}^{G}U_{gt}$, the planner allocates more unemployed workers to more effective labor markets, that is, with more vacancies and higher matching efficiency until their marginal contribution to the hires is equalized across markets. The main difference from Şahin et al. (2014) is that equilibrium condition (1) does not have an analytical formula like Cobb-Douglas specification. In Section 3.2, I explain how to compute the optimal allocation of unemployed workers numerically.
Using optimal allocation of unemployed workers $U_{gt}^{*}$, the aggregate actual and optimal number of new hires can be expressed as
$H_{t}$ $=\sum_{g=1}^{G}m(A_{gt}U_{gt},V_{gt}),$
$H_{t}^{*}$ $=\sum_{g=1}^{G}m(A_{gt}U_{gt}^{*},V_{gt}).$
@@ 3.2 Mismatch
$\max_{U_{1t},\cdots,U_{Gt}}$ $\sum_{g=1}^{G}m_{gt}\left(A_{gt}U_{gt},V_{gt}\right),$ (3)
subject to $\text{subject to }\frac{\partial m}{\partial U}\left(A_{1t}U_{1t}^{*},V_{1t}\right)$ $=\cdots=\frac{\partial m}{\partial U}\left(A_{Gt}U_{Gt}^{*},V_{Gt}\right),$
$U_{t}$ $=\sum_{g=1}^{G}U_{gt}.$
@@ 4 Data
First, I use the Report on Employment Service (Shokugyo Antei Gyomu Tokei) for the month-level aggregate data from January 1972 to April 2024. These datasets include the number of job openings, job seekers, and successful job placements, primarily sourced from the Ministry of Health, Labour and Welfare (MHLW) of Japan, which publishes monthly reports and statistical data on the Public Employment Security Office, commonly known as Hello Work. Hello Work is a government-operated institution in Japan that provides job seekers with employment counseling, job placement services, and vocational training, playing a critical role in Japan’s labor market. The data is often used as in Kano and Ohta (2005), Kambayashi and Ueno (2006), Sasaki (2008), and Higashi (2018) estimating the traditional Cobb-Douglas matching function. Using up-to-date Hello Work data, Kawata and Sato (2021) construct a simple framework to quantitatively measure the impacts of an economic shock of COVID-19 on unemployed workers’ welfare in their companion project. [^8] https://www.crepe.e.u-tokyo.ac.jp/material/crepecl12.html: Accessed 2024 June 6. The period for my dataset is selected to ensure the longest consistent timeframe available at the time of writing this paper. In Appendix A, we provide additional analysis using the year-level aggregate data in more extended periods, available from January 1963 to April 2024, instead of month-level one.
@@ 5.1 Aggregate trends in 1972-2023
Figures 2 (a)-(d) provide monthly data patterns of unemployed individuals, vacancies, labor market tightness ($V/U$), hires, the ($U$,$V$) relationship, and job and worker finding rates ($M/U$ and $M/V$) in aggregate labor markets. [^9] Worker finding rates are also known as vacancy yields. Davis et al. (2013) define vacancy yields in month $t$ as the number of hires in month $t\left(h_{t}\right)$ divided by the number of vacancies in the previous month $\left(v_{t-1}\right)$ instead of the same month $\left(v_{t}\right)$, the two series are very similar. In summary, the numbers of unemployed individuals and vacancies increase with fluctuations, corresponding with market tightness, while the number of hires shows a noticeable decline around the late 1970s, peaks and troughs from the mid-1980s to the late 1990s, peaks around mid-2000s and 2010, and a sharp decline towards recent years. Although the ($U$,$V$) relationship does not corresponds with Beveridge curve because these numbers are not divided by the corresponding number of labor force, the prominent rightward shifts are consistent with the U.S. studies (Elsby et al. 2015).
Figures 2 (e) and (f) present the estimation results of the matching function along with matching efficiency and elasticities ($\frac{d\ln M}{d\ln AU}$ and $\frac{d\ln M}{d\ln V}$). Notably, matching efficiency (normalized to 1972) shows a declining trend with notable fluctuations, which is consistent with the downward trends of job and worker finding rates. In particular, the significant decline after 2015 is remarkable. This seems due to an increase in matching opportunities outside of the government-operated platform, which is discussed in Otani (2024) using proprietary on-the-job search platform data.
@@ Appendix A Appendix
A.1 Year-level trends in 1966-2023

### v4

@@ Abstract
Keywords: matching efficiency, matching elasticity, matching function, mismatch
@@ 1.1 Related Literature
This paper contributes to the three strands of literature. First, I examine the trend of matching efficiency in Japanese labor markets via Hello Work nonparametrically using a novel approach (Lange and Papageorgiou 2020), which shows how to nonparametrically identify the matching function and estimate the matching function allowing for unobserved matching efficacy, without imposing the usual independence assumption between matching efficiency and search on either side of the labor market, allowing for multiple types of jobseekers. Lange and Papageorgiou (2020) highlight positive correlations between efficiency and market structure such as tightness and so on, which induces a positive bias in the estimates of the vacancy elasticity whenever unobserved matching efficacy is not controlled for, as is the case in the traditional Cobb-Douglas matching function with constant elasticity parameters. [^1] Brancaccio et al. (2020a, 2023) apply the method to estimate the matching function in exporter-ship transactions in the global bulk shipping market. Brancaccio et al. (2020b) summarize practical issues. Implementing their approach, I add updated results from the more flexible approach to the existing findings reported in Kano and Ohta (2005), Kambayashi and Ueno (2006), Sasaki (2008), and Higashi (2018) using the traditional Cobb-Douglas matching function with geographical and occupational category fixed effects to capture geographical and occupational heterogeneity. Table 1 summarizes the previous findings. My findings are also useful for comparison with other countries’ results reported in Bernstein et al. (2022) and Petrongolo and Pissarides (2001).
@@ 2.1 Nonparametric aggregate matching function
I identify the matching function as well as unobserved, time-varying matching efficiency, $A.$ First, I assume that $V$ and $A$ are independent conditional on $U$, that is, $A\perp V\mid U$. Second, I assume that the matching function $m(AU,V):\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}$ has constant returns to scale (CRS). [^5] To follow the original model of Matzkin (2003), we can rewrite $H=m(AU,V)$ into $H/U=m(A,V/U)$ under CRS. Here, $H/U$ and $V/U$ are known as a job-finding rate and market tightness. It is known that increasing returns to scale in the matching process can create multiple equilibria and the subsequent literature has found scant evidence for increasing returns, as in (Petrongolo and Pissarides 2001). Then, by applying nonparametric identification results of Matzkin (2003), Proposition 1 of Lange and Papageorgiou (2020) proves that $G(H,U,V)$ identifies $F(A,U)$ and $m(AU,V):\mathbb{R}_{+}^{2}\rightarrow\mathbb{R}_{+}$ up to a normalization of $A$ at one point denoted as $A_{0}$ of the support of $(A,U,V)$. [^6] In Appendix, I report finite sample performance and methodological extension with Monte Carlo simulation. Based on simulation results with sample size $T=50$, the sample size in this paper is enough to recover matching efficiency well. The code is available on the author’s Github. Also, see Brancaccio et al. (2020b) for practical issues. The approach is used to estimate a matching function in a trade model (Brancaccio et al. 2020a, 2023).
@@ 2.2 Nonparametric mismatch
I construct a nonparametric version of a mismatch index which is a measure of the fraction of hires foregone due to mismatch. Suppose there are $L$ local markets in time $t$ with a given number of vacancies, unemployed workers, and aggregate matching efficiencies in the economy. The number of hires $H_{\ell t}$ in market $g$ in time $t$ is determined by a matching function $m(A_{\ell t}U_{\ell t},V_{\ell t})$. Note that, unlike Şahin et al. (2014), I allow any specification of $m$ satisfying CRS and independence of $A_{\ell t}$ and $U_{\ell t}$ conditional on $V_{\ell t}$. Given matching efficiency $A_{\ell t}$ and the number of vacancies $V_{\ell t}$, a social planner maximizes the total hires in the economy by distributing a given number of unemployed workers to each labor market.
@@
$\frac{\partial m}{\partial U}\left(A_{1t}U_{1t}^{*},V_{1t}\right)=\cdots=\frac{\partial m}{\partial U}\left(A_{Lt}U_{Lt}^{*},V_{Lt}\right)$ (1)
where * denotes the planner’s allocation. The equilibrium condition means that keeping total unemployed workers $U_{t}=\sum_{\ell=1}^{L}U_{\ell t}$, the planner allocates more unemployed workers to more effective labor markets, that is, with more vacancies and higher matching efficiency until their marginal contribution to the hires is equalized across markets. The main difference from Şahin et al. (2014) is that equilibrium condition (1) does not have an analytical formula like Cobb-Douglas specification. In Section 3.2, I explain how to compute the optimal allocation of unemployed workers numerically.
Using optimal allocation of unemployed workers $U_{\ell t}^{*}$, the aggregate actual and optimal number of new hires can be expressed as
$H_{t}$ $=\sum_{\ell=1}^{L}m(A_{\ell t}U_{\ell t},V_{\ell t}),$
$H_{t}^{*}$ $=\sum_{\ell=1}^{L}m(A_{\ell t}U_{\ell t}^{*},V_{\ell t}).$
@@ 3.2 Mismatch
$\max_{U_{1t},\cdots,U_{Lt}}$ $\sum_{\ell=1}^{L}m\left(A_{\ell t}U_{\ell t},V_{\ell t}\right),$ (3)
subject to $\text{subject to }\frac{\partial m}{\partial U}\left(A_{1t}U_{1t}^{*},V_{1t}\right)$ $=\cdots=\frac{\partial m}{\partial U}\left(A_{Lt}U_{Lt}^{*},V_{Lt}\right),$
$U_{t}$ $=\sum_{\ell=1}^{L}U_{\ell t}.$
@@ 5 Data
First, I use the Report on Employment Service (Shokugyo Antei Gyomu Tokei) for the month-level aggregate data from January 1972 to April 2024. These datasets include the number of job openings, job seekers, and successful job placements, primarily sourced from the Ministry of Health, Labour and Welfare (MHLW) of Japan, which publishes monthly reports and statistical data on the Public Employment Security Office, commonly known as Hello Work. Hello Work is a government-operated institution in Japan that provides job seekers with employment counseling, job placement services, and vocational training, playing a critical role in Japan’s labor market. The data is often used as in Kano and Ohta (2005), Kambayashi and Ueno (2006), Sasaki (2008), Kambayashi (2013), and Higashi (2018) estimating the traditional Cobb-Douglas matching function. Using up-to-date Hello Work data, Kawata and Sato (2021) construct a simple framework to quantitatively measure the impacts of an economic shock of COVID-19 on unemployed workers’ welfare in their companion project. [^8] https://www.crepe.e.u-tokyo.ac.jp/material/crepecl12.html: Accessed 2024 June 6. The period for my dataset is selected to ensure the longest consistent timeframe available at the time of writing this paper. In Appendix A, we provide additional analysis using the year-level aggregate data in more extended periods, available from January 1963 to April 2024, instead of month-level one.
@@ 6.1 Aggregate trends in 1972-2023
Figures 2 (a)-(d) provide monthly data patterns of unemployed individuals, vacancies, labor market tightness ($V/U$), hires, the ($U$,$V$) relationship, and job and worker finding rates ($H/U$ and $H/V$) in aggregate labor markets. [^9] Worker finding rates are also known as vacancy yields. Davis et al. (2013) define vacancy yields in month $t$ as the number of hires in month $t\left(h_{t}\right)$ divided by the number of vacancies in the previous month $\left(v_{t-1}\right)$ instead of the same month $\left(v_{t}\right)$, the two series are very similar. In summary, the numbers of unemployed individuals and vacancies increase with fluctuations, corresponding with market tightness, while the number of hires shows a noticeable decline around the late 1970s, peaks and troughs from the mid-1980s to the late 1990s, peaks around mid-2000s and 2010, and a sharp decline towards recent years. Although the ($U$,$V$) relationship does not corresponds with Beveridge curve because these numbers are not divided by the corresponding number of labor force, the prominent rightward shifts are consistent with the U.S. studies (Elsby et al. 2015).
Figures 2 (e) and (f) present the estimation results of the matching function along with matching efficiency and elasticities ($\frac{d\ln m}{d\ln AU}$ and $\frac{d\ln m}{d\ln V}$). Notably, matching efficiency (normalized to 1972) shows a declining trend with notable fluctuations, which is consistent with the downward trends of job and worker finding rates. In particular, the significant decline after 2015 is remarkable. This seems due to an increase in matching opportunities outside of the government-operated platform, which is discussed in Otani (2024) using proprietary on-the-job search platform data and Kanayama and Otani (2024) using spot gig work platform data.
@@ Appendix A Appendix
A.1 Simulation results
  • How large is the bias of matching efficiency from the Cobb Douglass specification?
  • How large is the bias of mismatch from the Cobb Douglass specification?
  • Mismatch literature:
    • mismatch Shimer (2007)
    • mismatch and large firms (Eeckhout 2018)
    • Classical Jackman and Roper (1987)
  • We want to check robustness to specification, stationarity, and endogeneity.
    • Restricting CRS class
    • robustness to violation of independence of $A$ and $U$ conditional on $V$
  • We want to confirm how large enough the sample size is.
  • We want to quantify the bias of the standard Cobb-Douglas specification
A.2 Year-level trends in 1966-2023